When 70.6 g of calcium and 41.3 g of nitrogen gas undergo a reaction that has a 81.4% yield, what mass of calcium nitride forms?
1
Expert's answer
2016-09-20T05:24:10-0400
Solution:
According to the chemical equation of reaction:

3Ca + N2 = Ca3N2
Ar(Ca) = 40g/mole. Here of n(Ca) =70.6/40=1.77(mol)
Mr(N2)=28. Here of n(N2) = 41.3/28=1.48(mol)
3moles of Ca react with 1mole of N2, than N2 nitrogen in an excess. All calculation we will provide with Ca.

3Ca + N2 = Ca3N2
70.6/(3*40) = X/148

Where X=m(Ca3N2)g. here of X = 70.6*148/120=80.07g and 87.07*81.4/100 = 70.88(g)81.4%
Answer: mass of calcium nitride = 70.88